# Cross-Domain Graph Data Scaling: A Showcase with Diffusion Models

### v1

@@ 4 Experiment
4.1 Main results
Baselines. We evaluate our model against three main groups of baselines. (1) Task-specific GNNs: For graph property prediction, we use GIN xu2018powerful ; for link prediction, we use GCN kipf2017semisupervised and NCN wang2024neural ; and for node classification, we use GCN kipf2017semisupervised . (2) Graph pre-training methods: These include AttrMask, CtxtPred, EdgePred, and InfoMax hu2019strategies , JOAO you2021graph , D-SLA kim2022graph , and GraphMAE hou2022graphmae . For each of these methods, we pre-train it on the same pre-training set as UniAug. While most of the pre-training graphs lack node features, we calculate the node degrees as the input. Each method consists of three pre-trained variants with different backbone GNNs, including GIN, GCN, and GAT. We note that all these methods require the downstream graphs to have the same node feature space as the pre-training data. Therefore, in the fine-tuning stage, we replace the node features of the downstream datasets with node degrees, evaluate all three variants, and report the highest performance for each method in each task. (3) Graph data augmentation (GDA) methods: For graph property prediction, we include S-Mixup ling2023graph , GraphAug luo2022automated , FLAG kong2022robust , GREA liu2022graph , and DCT liu2024data ; for link prediction, we include CFLP zhao2022learning ; and for node classification on heterophilic graphs, we include Half-Hop azabou2023half . The GDA methods are implemented based on chosen task-specific GNNs.
Table 2: Mean and standard deviation of accuracy (%) with $10$-fold cross-validation on graph classification. The best result is bold. The highlighted results indicate negative transfer for pre-training methods compared to GIN. The last column is the average rank of each method.
DD Enzymes Proteins NCI1 IMDB-B IMDB-M Reddit-B Reddit-12K Collab A.R.
@@ 4.1 Main results
Graph property prediction. We employ graph label guidance for UniAug throughout the graph-level tasks by training a $2$-layer MLP as the guidance head on the graph labels in the training set. In the augmentation stage, we generate multiple graphs per training sample, and the generated graphs are then fed into the baseline GIN. We present the results of molecule regression in Table 3 and graph classification in Table 2. Three key observations emerge from the analysis: (1) Existing pre-training methods show negative transfer compared to GIN. Some special cases are the Enzymes and molecule regression datasets, where all pre-training methods fail to yield satisfactory results. In these datasets, the features are one of the driving components for graph property prediction, while the pre-training methods fail to encode such information due to incompatibility with the feature dimension. This reveals one critical drawback of the pre-training methods: their inability to handle feature heterogeneity. (2) GDA methods yield inconsistent results across different datasets. While these methods enhance performance in some datasets, they cause performance declines in others. This variability is directly reflected in the average rank, where some of them even fall behind the GIN. (3) Unlike the pre-training methods and GDA methods, UniAug shows consistent performance improvements against GIN with a large margin. In the molecule regression tasks, UniAug effectively compensates for the absence of bond features and achieves performance comparable to DCT, which is a data augmentation method pre-trained on in-domain molecule graphs. These findings affirm that the pre-training and structure augmentation paradigm of UniAug effectively benefits the downstream datasets at the graph level.
@@
Link prediction. We choose three guidance objectives for UniAug, including node degree, CN, and link prediction objective, as described in Section 3.3. For each objective, we train an MLP to provide guidance information. We then augment the graph structure by generating a synthetic graph and preserving the original training edges, ensuring that the augmented graph does not remove any existing edges. The augmented graph is then fed into a GCN for link prediction. We summarize the results in Table 4, which show similar patterns to those observed in graph property prediction: (1) Existing pre-training methods provide negative transfer, especially on datasets with node features. (2) GDA method CFLP leads to performance drops on the datasets without features and also suffers from high computation complexity during preprocessing. (3) UniAug enhances performance across all tested datasets. In addition, we employ UniAug to NCN wang2024neural , one of the state-of-the-art methods for link prediction. The results demonstrate consistent performance boosts from UniAug when we apply NCN as the backbone. The structure augmentation paradigm of UniAug allows plug-and-play applications to any downstream-specific models, showcasing its adaptability and effectiveness. In addition, we study the effects of three guidance objectives. More details can be found in Appendix D.2.
Table 4: Mean and standard deviation across $10$ runs on link prediction. Results are scaled ×$\times 100$. The last two methods are based on NCN, while the rest are GCN-based. The best result is bold for two backbones, respectively. The highlighted results indicate negative transfer for pre-training methods compared to GCN. The last column is the average rank of each GCN-based method.
Cora Citeseer Pubmed Power Yeast Erdos Flickr A.R.
MRR MRR MRR Hits@10 Hits@10 Hits@10 Hits@10
Cornell Wisconsin Texas Actor Chameleon* Squirrel* A.R.
  • *Chameleon and Squirrel are filtered to remove duplicated nodes platonov2023a .
@@
Table 6: Results of node classification on homophily graphs. Results are scaled ×$\times 100$.
@@
Node classification. To demonstrate the effectiveness of UniAug in node-level tasks, we transform the node classification into subgraph classification. Specifically, we extract the aggregation tree of each node, i.e., $2$-hop subgraph for a $2$-layer GCN, and label the subgraph with the center node. We then adopt a strategy similar to graph classification and train a $2$-layer classifier as a guidance head. Inspired by the success of structure augmentation on heterophilic graphs bi2022make ; azabou2023half , we evaluate UniAug on $6$ heterophilic datasets. We observe phenomena similar to those seen in graph property prediction and link prediction in Table 5. One thing to mention is the combination of UniAug and Half-Hop. Half-Hop offers performance improvements in four out of six datasets via data augmentation, and combining it with UniAug yields even higher results. This highlights the flexibility of UniAug and opens up possibilities for further exploration of its use cases. Given the impressive results of UniAug on heterophilic graphs, we anticipate it will also help to balance the performance disparities among nodes with different homophily ratios on homophilic graphs mao2024demystifying . We split the nodes into five groups according to their homophily ratios and calculate the standard deviation (SD) across groups. As shown in Table 6, UniAug matches the performance of vanilla GCN and also reduces the performance discrepancies corresponding to SD.
@@
[…]
4.2 Scaling behavior of UniAug
In light of the neural scaling law kaplan2020scaling ; hoffmann2022training ; abnar2022exploring ; zhai2022scaling ; liu2024neural , we expect UniAug to benefit from an increased volume of data and more compute budget. In this subsection, we investigate the scaling behavior of UniAug in terms of data scale and amount of compute for pre-training.
Data scale During the data collection process, we prepare three versions of the training data with increasing magnitude. We first sample $10$ graphs per category from the Network Repository rossi2015network to build a SMALL collection. Next, we gather all the graphs from the Network Repository and filter out large-scale graphs and outliers for a FULL collection. In addition, we add a $1000$ graphs subset of the GitHub Star dataset from TUDataset morris2020tudataset to enlarge the coverage of diverse patterns and form an EXTRA collection. We pre-train three versions of UniAug respectively on the three collections and evaluate them on graph classification and link prediction. As shown in Fig. 3, we observe a clear trend of increase in performance as we enlarge the scale of pre-training data. This paves the way to scale up UniAug to even more pre-training graphs.
Amount of compute We sought to understand how effectively our diffusion model can learn data patterns as we continue to train it. To this end, we checkpointed UniAug every 2,000 epochs ($5\times 10^{-3}$ PF-days) while training on the EXTRA collection, and then applied it to graph classification and link prediction tasks. The results are illustrated in Fig. 4. We observe that downstream performance generally improves with prolonged training, while the trend slows down for some datasets when we reach 8,000 epochs. We take the checkpoint at the 10,000th epoch for evaluations. Given the scaling behavior observed, we anticipate UniAug to become even more effective with additional resources.
@@ 5 Conclusion and Discussion
In this work, we propose a graph structure augmentation pipeline UniAug to leverage the increasing scale of graph data. We collect thousands of graphs from various domains and pre-train a self-conditioned discrete diffusion model on them. In the downstream stage, we augment the graphs by preserving the original node features and generating synthetic structures. We apply UniAug to node-, link-, and graph-level tasks and achieve consistent performance gain. We have successfully developed a showcase that benefits from cross-domain graph data scaling using diffusion models.
@@ Appendix A Derivation of Diffusion Process
$\begin{split}q\left(\mathbf{A}^{t}\mid\mathbf{A}^{t-1}\right)&=\operatorname{Bernoulli}\left(\mathbf{A}^{t};\alpha^{t}\mathbf{A}^{t-1}+\left(1-\alpha^{t}\right)\pi\right),\\ q\left(\mathbf{A}^{t}\mid\mathbf{A}^{0}\right)&=\operatorname{Bernoulli}\left(\mathbf{A}^{t};\bar{\alpha}^{t}\mathbf{A}^{0}+\left(1-\bar{\alpha}^{t}\right)\pi\right),\\ q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t},\mathbf{A}^{0}\right)&=\frac{q\left(\mathbf{A}^{t}\mid\mathbf{A}^{t-1}\right)q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{0}\right)}{q\left(\mathbf{A}^{t}\mid\mathbf{A}^{0}\right)},\end{split}$ (8)

### v3

@@ 4 Experiment
Table 2: Mean and standard deviation of accuracy (%) with $10$-fold cross-validation on graph classification. The best result is bold. The highlighted results indicate negative transfer for pre-training methods compared to GIN. The last column is the average rank.
$D$ Enzymes Proteins NCI1 IMDB-B IMDB-M Reddit-B Reddit-12K Collab A.R.
@@
4.1 Main results
Baselines.
We evaluate our model against three main groups of baselines. (1) Task-specific GNNs: For graph property prediction, we use GIN (xu2018powerful, ); for link prediction, we use GCN (kipf2017semisupervised, ) and NCN (wang2024neural, ); and for node classification, we use GCN (kipf2017semisupervised, ). (2) Graph pre-training methods: These include AttrMask, CtxtPred, EdgePred, and InfoMax (hu2019strategies, ), JOAO (you2021graph, ), D-SLA (kim2022graph, ), and GraphMAE (hou2022graphmae, ). For each of these methods, we pre-train it on the same pre-training set as UniAug. While most of the pre-training graphs lack node features, we calculate the node degrees as the input. Each method consists of three pre-trained variants with different backbone GNNs, including GIN, GCN, and GAT. (3) Self-supervised methods: These include MVGRL (hassani2020contrastive, ), GRACE (zhu2020deep, ), BGRL thakoor2022largescale and GraphMAE (hou2022graphmae, ). For the self-supervised methods, we extract the node embeddings and feed them into downstream specific heads. (4) Graph data augmentation (GDA) methods: For graph property prediction, we include S-Mixup (ling2023graph, ), GraphAug (luo2022automated, ), FLAG (kong2022robust, ), GREA (liu2022graph, ), and DCT (liu2024data, ); for link prediction, we include CFLP (zhao2022learning, ); and for node classification on heterophilic graphs, we include Half-Hop (azabou2023half, ). The GDA methods are implemented based on chosen task-specific GNNs.
Table 3: Mean and standard deviation across $10$ runs on link prediction. Results are scaled $\times 100$. The last two methods are based on NCN, while the rest are GCN-based. The best result is bold for two backbones, respectively. The last column is the average rank of each GCN-based method.
Cora Citeseer Pubmed Power Yeast Erdos Flickr A.R.
MRR MRR MRR Hits@10 Hits@10 Hits@10 Hits@10
GCN 30.26 ± 4.80 50.57 ± 7.91 16.38 ± 1.30 30.61 ± 4.07 24.71 ± 4.92 35.71 ± 2.65 8.10 ± 2.58 5.14
Self-supervised MVGRL 29.13 ± 3.90 51.32 ± 4.12 15.21 ± 2.35 31.71 ± 3.78 23.74 ± 5.74 36.21 ± 2.81 8.42 ± 2.18 5.29
GRACE 31.77 ± 4.31 49.13 ± 3.95 16.88 ± 1.74 28.21 ± 5.04 23.96 ± 4.31 33.90 ± 2.12 9.87 ± 0.98 5.14
BGRL 33.59 ± 2.14 51.91 ± 5.01 16.93 ± 2.03 33.71 ± 3.21 25.91 ± 3.12 37.95 ± 1.73 8.52 ± 1.85 3.00
GraphMAE 32.98 ± 5.01 52.71 ± 5.39 18.83 ± 1.30 32.81 ± 2.12 26.51 ± 2.92 35.63 ± 3.61 7.01 ± 3.86 3.43
GDA CFLP 33.62 ± 6.44 55.20 ± 4.16 17.01 ± 2.75 16.02 ± 8.31 24.23 ± 5.23 28.74 ± 2.38 0.00 ± 0.00 4.57
UniAug - GCN 35.36 ± 7.88 54.66 ± 4.55 17.28 ± 1.89 34.36 ± 1.68 27.52 ± 4.80 39.67 ± 4.51 9.46 ± 1.18 1.43
NCN-based NCN 31.72 ± 4.48 58.03 ± 3.45 38.26 ± 2.56 27.36 ± 5.00 39.85 ± 5.07 36.81 ± 3.29 8.33 ± 0.92
Table 4: Mean and standard deviation of accuracy (%) across $10$ splits on node classification of heterophilic graphs. The best result is bold. The last column is the average rank of each method.
Cornell Wisconsin Texas Actor Chameleon* Squirrel* A.R.
GCN 59.41 ± 6.03 51.68 ± 4.34 63.78 ± 4.80 30.58 ± 1.29 40.94 ± 3.91 39.11 ± 1.74 4.50
Self-supervised MVGRL 56.19 ± 2.42 50.64 ± 5.89 61.70 ± 3.94 31.37 ± 0.83 32.34 ± 2.11 35.32 ± 1.32 6.83
GRACE 56.39 ± 2.11 53.83 ± 3.56 63.54 ± 2.57 28.14 ± 0.81 35.71 ± 1.95 33.65 ± 2.51 7.00
BGRL 56.67 ± 2.13 59.80 ± 4.08 65.78 ± 2.66 29.80 ± 0.31 37.01 ± 2.89 34.77 ± 2.01 5.33
GraphMAE 57.31 ± 2.11 58.27 ± 2.91 58.34 ± 3.57 28.97 ± 0.27 36.75 ± 1.78 39.13 ± 2.01 5.50
GDA Half-Hop 62.46 ± 7.58 76.47 ± 2.61 72.35 ± 4.27 33.95 ± 0.68 38.59 ± 2.89 37.34 ± 2.18 3.17
UniAug + Half-Hop 72.43 ± 5.81 79.61 ± 5.56 77.03 ± 4.27 34.97 ± 0.55 41.94 ± 2.77 38.79 ± 2.61 1.67
@@ 4.1 Main results
We employ graph label guidance for UniAug throughout the graph-level tasks by training a $2$-layer MLP as the guidance head on the graph labels in the training set. In the augmentation stage, we generate multiple graphs per training sample, and the generated graphs are then fed into the baseline GIN. We present the results of molecule regression in Table 5 and graph classification in Table 2. Three key observations emerge from the analysis: (1) Existing pre-training methods show negative transfer compared to GIN. Some special cases are the Enzymes and molecule regression datasets, where all pre-training methods fail to yield satisfactory results. In these datasets, the features are one of the driving components for graph property prediction, while the pre-training methods fail to encode such information due to incompatibility with the feature dimension. This reveals one critical drawback of the pre-training methods: their inability to handle feature heterogeneity. (2) GDA methods yield inconsistent results across different datasets. While these methods enhance performance in some datasets, they cause performance declines in others. This variability is directly reflected in the average rank, where some of them even fall behind the GIN. (3) Unlike the pre-training methods and GDA methods, UniAug shows consistent performance improvements against GIN with a large margin. In the molecule regression tasks, UniAug effectively compensates for the absence of bond features and achieves performance comparable to DCT, which is a data augmentation method pre-trained on in-domain molecule graphs. We also adapt the semi-supervised (you2020l2, ) and self-supervised (Sun2020InfoGraph, ) setting for the baselines for a comprehensive benchmark in Appendix D.1 Table 13, where we observe that UniAug presents consistently satisfactory performance according to the average rank, matching or outperforming the best baseline. These findings affirm that the pre-training and structure augmentation paradigm of UniAug effectively benefits the downstream datasets at the graph level.
Link prediction.
@@
We choose three guidance objectives for UniAug, including node degree, CN, and link prediction objective, as described in Section 3.3. For each objective, we train an MLP to provide guidance information. We then augment the graph structure by generating a synthetic graph and preserving the original training edges, ensuring that the augmented graph does not remove any existing edges. The augmented graph is then fed into a GCN for link prediction. We summarize the results in Table 3, where we observe that (1) GDA method CFLP leads to performance drops on the datasets without features and also suffers from high computation complexity during preprocessing. (2) UniAug enhances performance across all tested datasets. In addition, we employ UniAug to NCN (wang2024neural, ), one of the state-of-the-art methods for link prediction. The results demonstrate consistent performance boosts from UniAug when we apply NCN as the backbone. The structure augmentation paradigm of UniAug allows plug-and-play applications to any downstream-specific models, showcasing its adaptability and effectiveness. Additional results, including the performance of pre-training baselines and the effects of three guidance objectives, can be found in Appendix D.2.
@@
Table 6: Results of node classification on homophily graphs. Results are scaled $\times 100$.
@@
Node classification.
To demonstrate the effectiveness of UniAug in node-level tasks, we transform the node classification into subgraph classification. Specifically, we extract the aggregation tree of each node, i.e., $2$-hop subgraph for a $2$-layer GCN, and label the subgraph with the center node. We then adopt a strategy similar to graph classification and train a $2$-layer classifier as a guidance head. Inspired by the success of structure augmentation on heterophilic graphs (bi2022make, ; azabou2023half, ), we evaluate UniAug on $6$ heterophilic datasets. We observe phenomena similar to those seen in graph- and link-level tasks in Table 4, with results of pre-training baselines in Appendix D.3. One thing to mention is the combination of UniAug and Half-Hop. Half-Hop offers performance improvements in four out of six datasets via data augmentation, and combining it with UniAug yields even higher results. This highlights the flexibility of UniAug and opens up possibilities for further exploration of its use cases. Given the impressive results of UniAug on heterophilic graphs, we anticipate it will also help to balance the performance disparities among nodes with different homophily ratios on homophilic graphs (mao2024demystifying, ). We split the nodes into five groups according to their homophily ratios and calculate the standard deviation (SD) across groups. As shown in Table 6, UniAug matches the performance of vanilla GCN and also reduces the performance discrepancies corresponding to SD.
4.2 Scaling behavior of UniAug
In light of the neural scaling law (kaplan2020scaling, ; hoffmann2022training, ; abnar2022exploring, ; zhai2022scaling, ; liu2024neural, ), we expect UniAug to benefit from an increased coverage of data and more compute budget. In this subsection, we investigate the scaling behavior of UniAug in terms of data scale and amount of compute for pre-training.
Data coverage
During the data collection process, we prepare three versions of the training data with increasing magnitude and growing coverage on the graph distribution. We first sample $10$ graphs per category from the Network Repository (rossi2015network, ) to build a SMALL collection. Next, we gather all the graphs from the Network Repository and filter out large-scale graphs and outliers for a FULL collection. In addition, we add a $1000$ graphs subset of the GitHub Star dataset from TUDataset (morris2020tudataset, ) to enlarge the coverage of diverse patterns and form an EXTRA collection. We pre-train three versions of UniAug respectively on the three collections and evaluate them on graph classification and link prediction. As shown in Fig. 3, we observe a clear trend of increase in performance as we enlarge the coverage of pre-training data. This paves the way to scale up UniAug to even more pre-training datasets with an expanding distribution of graphs. Additional results on scaling effects with respect to the proportion of the full data can be found in Appendix D.4.
Amount of compute
We sought to understand how effectively our diffusion model can learn data patterns as we continue to train it. To this end, we checkpointed UniAug every 2,000 epochs ($5\times 10^{-3}$ PF-days) while training on the EXTRA collection, and then applied it to graph classification and link prediction tasks. The results are illustrated in Fig. 4. We observe that downstream performance generally improves with prolonged training, while the trend slows down for some datasets when we reach 8,000 epochs. We take the checkpoint at the 10,000th epoch for evaluations. Given the scaling behavior observed, we anticipate UniAug to become even more effective with additional resources.
@@ 4.2 Scaling behavior of UniAug
Figure 4: Effects of pre-training amount of compute on graph classification (left) and link prediction (right), where one $\text{PF-days}=10^{15}\times 24\times 3600=8.64\times 10^{19}$ floating point operations.
@@ 5 Conclusion and Discussion
In this work, we propose a graph structure augmentor UniAug to leverage the increasing scale of graph data. We collect thousands of graphs from various domains and pre-train a self-conditioned discrete diffusion model on them. In the downstream stage, we augment the graphs by preserving the original node features and generating synthetic structures. We apply UniAug to node-, link-, and graph-level tasks and achieve consistent performance gain. We have successfully developed a showcase that benefits from cross-domain graph data scaling using diffusion models.
@@
6 Impact Statements
7 Acknowledgment
W.T. and J.T. are supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IIS 2504089, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, the Michigan Department of Agriculture and Rural Development, US Dept of Commerce, Amazon Faculty Award, Meta, NVIDIA, Microsoft and SNAP. W.T. and Y.X. are supported by the National Institutes of Health (NIH) under grant numbers U01DE033330, R01HL166508, R01DE026728 and the NSF under grant numbers OISE2434687 and IOS2107215.
8 Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates (“JP Morgan”), and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
@@ Appendix A Derivation of Diffusion Process
$\begin{split}q\left(\mathbf{A}^{t}\mid\mathbf{A}^{t-1}\right)&=\operatorname{Bernoulli}\left(\mathbf{A}^{t};\alpha^{t}\mathbf{A}^{t-1}+\left(1-\alpha^{t}\right)\pi\right),\\ q\left(\mathbf{A}^{t}\mid\mathbf{A}^{0}\right)&=\operatorname{Bernoulli}\left(\mathbf{A}^{t};\bar{\alpha}^{t}\mathbf{A}^{0}+\left(1-\bar{\alpha}^{t}\right)\pi\right),\\ q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t},\mathbf{A}^{0}\right)&=\frac{q\left(\mathbf{A}^{t}\mid\mathbf{A}^{t-1}\right)q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{0}\right)}{q\left(\mathbf{A}^{t}\mid\mathbf{A}^{0}\right)},\end{split}$
@@ D.1 Graph property prediction
In Section 4.1, we aim to benchmark the capability of cross-domain pre-training of different methods on the same set of pre-training graphs. While the pre-training graphs contain vastly different features, we have to align the feature space to allow pre-training for the baseline methods. There are two ways to tackle the feature heterogeneity issues in the existing literature. One line of them utilizes LLMs to align text-space graphs [96], which is not applicable to broader classes of graphs. Other works, like GCOPE [97], perform dimension reduction to align the feature dimension of different graphs. We emphasize that dimension reduction methods fail to deal with extreme cases like missing features. This phenomenon is pretty common in real life, as a large proportion of the graphs in the Network Repository do not have corresponding features. Therefore, we simply use the node degrees as the features in Section 4.1.
We understand that removing the node features may result in a performance drop for the baseline methods. Note that most of the baselines follow the pre-training paradigm of [22] with domain-specific model designs for chemistry and biology datasets, and thus cannot be directly applied to the chosen graph classification datasets. Therefore, we adapt the semi-supervised [76] and self-supervised [77] setting for the baselines for a comprehensive benchmark. The semi-supervised setting involves pre-training with all data of that specific dataset and finetuning the training set of each split. Meanwhile, baselines of the self-supervised setting pre-train on the whole dataset and then classify the learned graph embeddings with a downstream SVM classifier. The results are summarized in Table 13, where the best and second-best results are highlighted in bold and italic, respectively. We observe that UniAug presents consistently satisfactory performance according to the average rank, matching or outperforming the best baseline.
Table 13: Mean and standard deviation of accuracy (%) with $10$-fold cross-validation on graph classification. The best and second-best results are highlighted in bold and italic. The last column is the average rank.
$D$ Proteins NCI1 IMDB-B IMDB-M Reddit-B Collab A.R.
Semi-supervised CtxtPred 74.66±0.51 70.23±0.63 73.00±0.30 88.66±0.95 73.69±0.37 6.80
InfoMax 75.78±0.34 72.27±0.40 74.86±0.26 88.66±0.95 73.76±0.29 5.60
GraphCL 76.17±1.37 74.17±0.34 74.63±0.25 89.11±0.19 74.23±0.21 4.60
JOAO 75.81±0.73 73.31±0.48 74.86±0.39 88.79±0.65 75.53±0.18 4.60
Self-supervised InfoGraph 74.44±0.31 76.20±1.06 73.03±0.87 49.69±0.53 82.50±1.42 70.65±1.13 5.17
GraphCL 74.39±0.45 77.87±0.41 71.14±0.44 48.58±0.67 89.53±0.84 71.36±1.15 4.50
JOAO 74.55±0.41 78.07±0.47 70.21±3.08 49.20±0.77 85.29±1.35 69.50±0.36 5.17
GraphMAE 75.30±0.39 80.40±0.30 75.52±0.66 51.63±0.52 88.01±0.19 80.32±0.46 2.17
UniAug 78.13±2.61 75.47±2.50 80.54±1.77 73.50±2.48 50.13±2.05 92.28±1.59 77.00±2.02 1.43
To demonstrate the effectiveness of UniAug in scenarios with limited labeled data, we perform 5-shot graph classification following [98]. The results are summarized in Table 14 These results show that UniAug achieves significant performance improvements over the self-supervised baselines, underscoring its robustness and adaptability in few-shot settings.
Table 14: Accuracy of 5-shot graph classification.
Proteins Enzymes
GIN 58.17 ± 8.58 20.34 ± 5.01
InfoGraph 54.12 ± 8.20 20.90 ± 3.32
GraphCL 56.38 ± 7.24 28.11 ± 4.00
JOAO 57.21 ± 6.91 35.31 ± 3.79
GraphMAE 60.03 ± 5.35 33.91 ± 6.58
UniAug 66.85 ± 4.71 48.37 ± 4.77
In addition, to showcase the flexibility of UniAug on the downstream backbone, we pick one of the SOTA method PIN [99] for graph classification and evaluate UniAug on the basis of it. The results are summarized in Table 15, where we see UniAug offers constant improvements over PIN.
Table 15: Accuracy of graph classification with PIN.
Proteins NCI1 IMDB-B
PIN 78.8 ± 4.4 85.1 ± 1.5 76.6 ± 2.9
UniAug - PIN 80.2 ± 2.8 86.5 ± 1.4 77.9 ± 1.8